Norsk Optronics, ALS, of Bergen, Norway, had a current ratio of 2 on June 30 of the current year. On that date, the company’s assets were:

Cash $ 62,000
Accounts receivable, net 440,000
Inventory 690,000
Prepaid expenses 11,000
Plant and equipment, net 1,790,000
Total assets $ 2,993,000

Required:

1. What was the company’s working capital on June 30?

2. What was the company’s acid-test ratio on June 30? (Round your answer to 2 decimal places.)

3. The company paid an account payable of $48,000 immediately after June 30.

a. What effect did this transaction have on working capital?

b. What effect did this transaction have on the current ratio? (Round your intermediate calculations to 1 decimal place.)
Answer #1
June 30
Cash 62000
Accounts receivables, net 440000
Inventory 690000
Prepaid expenses 11000
Current assets 1203000
1.
Current ratio = Current assets / Current liabilities
2 = 1203000 / Current liabilities
Current liabilities = 1203000 / 2 601500
Working capital = Current assets - Current liabilities = 1203000 - 601500 601500
2.
Acid-test ratio = ( Current assets - Inventory - Prepaid expenses ) / Current liabilities = ( 1203000 - 690000 - 11000 ) / 601500 0.83
3.
After payment of 48000 to accounts payables, both current assets and current liabilities will decrease by 48000
Current assets after payment = 1203000 - 48000 1155000
Current liabilities after payment = 601500 - 48000 553500
a. Working capital after payment = Current assets after payment - Current liabilities after payment = 1155000 - 553500 601500
Answer : Working capital would not be affected
b. Current ratio after payment = Current assets after payment / Current liabilities after payment = 1155000 / 553500 2.1
Answer : Current ratio would increase